Find the potential difference across a resistor of resistance 10 ohm, if 10 to the power 20 electrons flow through it in one second .

Grade:10

4 Answers

Utsav Basu
6 years ago
1020 electron pass through the resistor in one second.
So, charge flowing through it = 1020 * 1.6 * 10-19 C = 16 C
Current = 16/1 A = 16 A
Now, using V = IR, we get,
Potential difference = 16 A * 10 Ω = 160 V
Gitanjali Rout
5 years ago
1020 electron pass through the resistor in one second.
So, charge flowing through it = 1020 * 1.6 * 10-19 C = 16 C
Current = 16/1 A = 16 A
Now, using V = IR, we get,
Potential difference = 16 A * 10 Ω = 160 V
Gitanjali Rout
5 years ago
So, charge flowing through it = 1020 * 1.6 * 10-19 C = 16 C
Current = 16/1 A = 16 A
Now, using V = IR, we get,
Potential difference = 16 A * 10 Ω = 160 V
Gitanjali Rout
5 years ago
So, charge flowing through it = 1020 * 1.6 * 10-19 C = 16 C
Current = 16/1 A = 16 A
Now, using V = IR, we get,
Potential difference = 16 A * 10 Ω = 160 V.